A race car accelerates uniformly from 18.5 m/s to 46.1 m/s in 2.47 seconds. Determine the acceleration of the car and the distance traveled.
Answer by stanbon(48558)

A race car accelerates uniformly from 18.5 m/s to 46.1 m/s in 2.47 seconds. Determine the acceleration of the car and the distance traveled.
------------------------------------
Average speed = (46.1-18.5)/2.47 = 11.174 mps
---------------------------
Distance = rate * time = 11.174*2.47 = 27.6 m
================================================
Cheers,
Stan H.